### v1

# AnTKV: Anchor Token-Aware Sub-Bit Vector Quantization for KV Cache in Large Language Models

Abstract
Quantization has emerged as an effective and lightweight solution to reduce the memory footprint of the KV cache in Large Language Models (LLMs). Nevertheless, minimizing the performance degradation caused by ultra-low-bit KV cache quantization remains a significant challenge. We observe that quantizing the KV cache of different tokens has varying impacts on the quality of attention outputs. To systematically investigate this phenomenon, we perform forward error propagation analysis on attention and propose the Anchor Score (AnS) that quantifies the sensitivity of each token’s KV cache to quantization-induced error. Our analysis reveals significant disparities in AnS across tokens, suggesting that preserving a small subset with full precision (FP16) of high-AnS tokens can greatly mitigate accuracy loss in aggressive quantization scenarios. Based on this insight, we introduce AnTKV, a novel framework that leverages Anchor Token-aware Vector Quantization to compress the KV cache. Furthermore, to support efficient deployment, we design and develop a triton kernel that is fully compatible with FlashAttention, enabling fast online Anchor Token selection. AnTKV enables LLaMA-3-8B to handle context lengths up to 840K tokens on a single 80GB A100 GPU, while achieving up to 3.5x higher decoding throughput compared to the FP16 baseline. Our experiment results demonstrate that AnTKV matches or outperforms prior works such as KIVI, SKVQ, KVQuant, and CQ under 4-bit settings. More importantly, AnTKV achieves significantly lower perplexity under ultra-low-bit quantization on Mistral-7B, with only 6.32 at 1-bit and 8.87 at 0.375-bit, compared to the FP16 baseline of 4.73.
[…]
1 Introduction
LLMs have attracted much attention due to their strong capability in natural language processing and other artificial intelligence tasks openai2023gpt4 ; guo2025deepseek ; team2023gemini ; tang2025the ; dong2025can . With the rapid development in recent years, LLM can manage increasingly larger context lengths ranging from hundreds of thousands to millions of tokens, and its broadening scope allows LLM to tackle more complex tasks zhao2023survey ; wang2025all ; das2025security ; wang2025agenttaxo . However, these models, which are typically built on decoder-based transformer architectures, generate tokens recursively during inference, thus leading to a large accumulation of the KV cache as the context length zhu2025oraclekv ; bai2024longbench ; liu2025comprehensive and batch size increases pope2022efficientlyscalingtransformerinference . The large memory footprint of the KV cache presents a significant challenge. For long-context inference, such as LLaMA-3’s 128K context length, the KV cache can consume several gigabytes of memory, nearly matching the model’s weights. This challenge is even greater for models like Gemini, which handle contexts up to a million tokens. Moreover, LLM-RAG systems yao2025cacheblend pre-generate and store KV caches. However, the excessive size of these caches presents a significant challenge for storage.
Figure 1: Self-attention output error when quantizing the KV cache of Mistral-7B to 1-bit.
Figure 2: The perplexity of Mistral-7B on the Wikitext2 under different bit-widths.
To address the growing KV cache, various optimization strategies have been proposed liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; liu2024retrievalattention ; zhang2024cam ; liu2024kivi ; dong2025can , with quantization emerging as a prominent solution. Quantization approaches fall into two main categories: scalar quantization (SQ) and vector quantization (VQ). SQ compresses the KV cache by mapping floating-point values to fixed low-bit representations liu2024kivi ; duanmu2024skvq ; hooper2024kvquant , but the minimum 1-bit requirement limits its compression to 1/16 on FP16. In contrast, VQ maps high-dimensional vectors to a finite codebook lingle2024transformervq ; zhang2024kv , capturing intra-vector correlations. This enables higher compression rates and better preserves representational fidelity, especially under ultra-low-bit regimes.
During the LLM inference, different tokens have different effects on the accuracy of LLMs, so it is unreasonable to treat all tokens equally in VQ methods. To illustrate this, we quantize the KV cache of Mistral-7B mistral to 1-bit using VQ, and Figure 1 shows the impact of KV quantization on the self-attention output error within layer 31. It reveals that even in the absence of tokens with exceptionally large quantization errors, a few tokens (Anchor Tokens) can still introduce substantial deviations in the output of the self-attention operator. Therefore, measuring the importance of tokens during inference can further explore the performance potential of VQ methods for compressing the KV cache.
To this end, we introduce Anchor Score (AnS), a metric derived from forward error analysis, to quantify each token’s impact on the output. Leveraging AnS, we propose AnTKV, a novel framework that employs anchor token-aware VQ for KV cache compression. AnTKV mitigates quantization-induced degradation through a two-stage design: offline token-aware centroids learning and online anchor token selection. In the offline stage, the framework performs the token-aware weighted k-means clustering to generate centroids. During the online stage, it computes the AnS for each request to identify anchor tokens that significantly contribute to output error and selectively retains their FP16. Moreover, to ensure efficiency, we expose key intermediate variables from FlashAttention dao2023flashattention2 and integrate a lightweight GPU kernel for AnS computation during inference, enabling efficient runtime anchor token selection.
To validate AnTKV’s effectiveness, we conduct extensive experiments across various quantization bit-widths on a range of LLMs, including LLaMA-2/3 llama2 ; llama3 and Mistral-7B mistral family. As depicted in Figure 2, AnTKV consistently achieves better performance across a wide spectrum of bit-widths, from 4-bit down to 0.375-bit. On Wikitext2 with Mistral-7B, AnTKV achieves a perplexity of 6.32 at 1-bit, a reduction of 0.93 compared to CQ (7.25) zhang2024kv , and substantial 9.04 over KVQuant-1% (15.36) hooper2024kvquant . Even at an aggressive 0.375-bit, AnTKV further lowers perplexity to 8.87, surpassing CQ (13.20) by 4.33. Enabled by the efficient implementation of AnS, AnTKV scales to extremely long contexts. On LLaMA-3-8B, it supports up to 840K context length under 0.375-bit on a single A100-80GB GPU. During decoding, AnTKV increases the maximum batch size by 3.3x and improves throughput by 3.4x at 1K context length compared to FP16. More experiments that demonstrate the performance advantages of the proposed AnTKV are in Section 4.
To summarize, we make the following contributions in this work.
  • We conduct an error analysis of the self-attention, and introduce AnS to quantify the importance of tokens. With AnS, we can identify anchor tokens that significantly impact the accuracy of LLMs. Based on AnS, we present a novel token-aware VQ method, namely AnTKV, to compress the KV cache in LLM inference.
  • To ensure efficiency, we design a custom GPU kernel for computing AnS in the online stage that is compatible with FlashAttention. With the high-performance implementation of AnS, AnTKV can support up to 840K context length on a single GPU.
  • We deploy the proposed AnTKV in LLaMA-2/3 and Mistral-7B family, numerical results show that AnTKV has a significant performance improvement under various quantization bit-widths.
2 Background and Related Works
2.1 Transformer and Attention
The transformer block has become the fundamental architecture of LLMs; it consists of a fully connected feed-forward layer and a self-attention layer. The self-attention layer enables the model to capture connections among tokens within the context and serves as the critical component of the transformer block. Specifically, it maps a query ($\bm{Q}$) and a set of KV pairs ($\bm{K}$ and $\bm{V}$) to an output, i.e.,
$\texttt{Attn}(\bm{Q},\bm{K},\bm{V})=\texttt{Softmax}\left(\frac{\bm{Q}\bm{K}^{T}}{\sqrt{d}}\right)\bm{V}\quad\text{with}\quad\bm{Q,K,V}\in\mathbb{R}^{n\times d},$
where $\texttt{Softmax}(\cdot)$ is the softmax operator. Since position information is crucial in LLMs, Rotary Position Embedding (RoPE) su2023roformerenhancedtransformerrotary is a widely used technique for encoding it into query and key vectors su2024roformer , which is denoted as $\widetilde{\bm{Q}}$ and $\widetilde{\bm{K}}$ in this paper. With RoPE, the attention score $\bm{A}$ is rewritten as $\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right)$, and its $L_{p}$ “entry-wise” matrix norm is defined as $\left(\sum\limits_{i,j}|\bm{A}_{i,j}|^{p}\right)^{1/p}$. Other notations used in this work, such as the Kronecker product, the Hadamard product, and the row-by-row vectorization are denoted as $\otimes$, $\odot$, and $\texttt{Vec}(\cdot)$, respectively.
2.2 Memory Constraints in LLM Inference
During the process of LLM inference, the key and value generated at prefill and each decoding step are stored to avoid redundant computation. The KV cache is repeatedly accessed in subsequent steps to compute attention over the full context seen so far, significantly accelerating autoregressive decoding. In a multi-turn conversation, the KV cache from previous turns can also be reused during the prefill stage to further reduce latency. The state-of-the-art open-source and closed-source LLMs, like LLaMA-3.1-8B llama3 and Gemini 1.5 Pro team2023gemini , now handle much longer context lengths, up to 128K and 2 million tokens, respectively. This increase in context length liu2025comprehensive ; wang2025agenttaxo causes the GPU memory used for the KV cache in LLM inference to even exceed the models’ size. The rapid growth of KV cache greatly limits the deployment of LLMs with long context. Quantizing the KV cache directly optimizes this by maintaining key-value information and cutting memory costs substantially, thus enhancing the efficiency of LLM inference.
2.3 Related Works
KV cache quantization A variety of KV cache quantization methods have been developed to address the memory bottleneck in long-context LLMs liu2024kivi ; duanmu2024skvq ; hooper2024kvquant ; zhang2024kv ; zhu2024sampleattentionnearlosslessaccelerationlong . KIVI liu2024kivi reduces quantization error via per-channel key quantization combined with a sliding window strategy to emphasize locally relevant tokens. SKVQ duanmu2024skvq further explores this direction by introducing channel reordering and clipping to improve low-bit quantization. In contrast, KVQuant hooper2024kvquant adopts a different design by quantizing pre-RoPE keys using non-uniform quantization and selectively preserving a small fraction of full-precision KV entries to mitigate accuracy loss. CQ zhang2024kv adopts a VQ-based approach, aiming to exploit cross-channel correlations to further compress the KV cache.
KV cache compression Beyond quantization, the field of LLMs is actively exporing several advanced methods for KV cache compression. Sparse attention are designed to achieve compression by selectively computing attention over a reduced set of tokens xiaoefficient ; chen2024prefixquant ; zhu2025oraclekv ; liu2025chunkkv ; liu2025llmsmaintainfundamentalabilities ; li2024snapkvllmknowslooking , However, a significant limitation lies in the inherent complexity of designing and optimizing their sparsity patterns. This often necessitates task-specific or data-driven heuristics, which can severely constrain their efficient generalization across diverse model architectures and datasets. Token merge, as a variant of sparse attention, merges the KV cache of similar tokens during inference to reduce memory usage zhang2024cam ; wang2024modeltellsmergeadaptive . Retrieval-based methods liu2024retrievalattention ; chenmagicpig ; zhang2024pqcache compress KV caches by retrieving relevant historical context for each query and are effective in tasks like RAG or document retrieval. However, they often lose substantial KV information, making them less suitable for general-purpose LLM inference.
Model Compression Many model compression techniques closely align with KV cache compression in both goals and underlying methodologies. GPTQ frantar2023gptqaccurateposttrainingquantization utilizes calibration set to reduce quantization induced degradation, while SmoothQuant xiao2024smoothquantaccurateefficientposttraining and AWQ lin2024awqactivationawareweightquantization minimize output error from the perspective of error propagation analysis. VQ-based methods such as VPTQ liu2024vptqextremelowbitvector and QUIP# tseng2024quipbetterllmquantization further enhance compression fidelity through residual VQ and Hadamard transform. Additionally, system-level frameworks like Atom zhao2024atomlowbitquantizationefficient and QServe lin2025qservew4a8kv4quantizationcodesign jointly optimize model and KV cache quantization, achieving end-to-end acceleration under low-bit quantization. Other works, such as Pruner-Zero dong2024prunerzeroevolvingsymbolicpruning and Parzc dong2024parzcparametriczerocostproxies , explore how to sparsify model weights while preserving model performance. Systems like FlashLLM xia2023flashllmenablingcosteffectivehighlyefficient and Spinfer fan2025spinfer leverage the model sparsity to accelerate inference.
3 Methodology
Previous works on quantifying the KV cache typically do not distinguish the importance of individual tokens, but assume that all tokens contribute equally to attention accuracy, or they adopt a simpler strategy of only preserving designated sink tokens. However, our 1-bit VQ experiment on Mistral-7B contradicts this, unveiling a previously unaddressed aspect in KV cache quantization: a small subset of tokens provokes disproportionately large perturbations in the self-attention outputs. This divergence reveals a new opportunity for KV cache quantization. Instead of uniform application, the precision of quantization should correspond to each token’s individual importance. Specifically, if we identify and preserve a subset of high-impact tokens in FP16, we can unlock a far better memory–accuracy trade-off.
Inspired by this, we present a novel token-aware VQ method for KV cache compression, namely, AnTKV. In AnTKV, we adopt a two-stage design: offline token-aware centroids learning (see Figure 3(a)) and online anchor token selection (see Figure 3(b)), the key of which lies in how to measure the importance of tokens. In the offline stage, we perform a weighted k-means clustering to produce centroids in VQ, and the gradient corresponding to each token is used as weights. In the online stage, to reduce the output error caused by quantization, we select tokens with both higher $\text{Err}_{\text{quant}}$ and AnS, i.e. anchor tokens, and keep them at FP16. Additionally, since applying RoPE disrupts the consistency of channel-wise magnitudes in $\bm{K}$ (see Appendix A for an illustrative example), which exhibits large inter-cluster distances and small intra-cluster variances, leading to large quantization error, the pre-RoPE strategy with the same as hooper2024kvquant is used in AnTKV.
Figure 3: The overview of AnTKV. (a) In the offline stage, we perform token-aware centroid learning to obtain the centroids. (b) In the online stage, we apply VQ using centroids and compute the AnS. Then identify the top-$k$ anchor tokens with the highest product of VQ error and AnS, and retain them in full-precision to preserve model performance.
3.1 Offline Token-Aware Centroids Learning
The offline stage aims to learn centroids by aggregating KV cache from general calibrations. Specifically, each row (per token) of $\bm{K}$ and $\bm{V}$ is divided into several sub-vectors, then using k-means to cluster them, and only the centroids are stored. To integrate the importance of tokens, we introduce the weighted k-means instead of classical k-means for the offline centroids learning.
From the perspective of accuracy, we hope that the errors caused by quantizing $\bm{K}$ and $\bm{V}$ will not affect the accuracy of LLMs as much as possible. Mathematically, it can be formulated as
$\begin{split}\min\limits_{\bm{C}^{k}\in\mathbb{R}^{c^{k}\times d}}|L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))|,\\ \end{split}$ (1)
where $c^{k}$ is the number of clusters, and $\bm{C}^{k}$ is the matrix consisting of centroids. By the first-order Taylor series expansion, we have
$\begin{split}L(\bm{K},\bm{V})-L(\texttt{vq}(\bm{K}),\texttt{vq}(\bm{V}))\approx\sum\limits_{j}\left(\langle\frac{\partial L}{\partial\bm{K}_{j,:}},\bm{K}_{j,:}-\texttt{vq}(\bm{K}_{j,:})\rangle+\langle\frac{\partial L}{\partial\bm{V}_{j,:}},\bm{V}_{j,:}-\texttt{vq}(\bm{V}_{j,:})\rangle\right),\end{split}$ (2)
where $\frac{\partial L}{\partial\bm{K}_{j,:}}$ and $\frac{\partial L}{\partial\bm{V}_{j,:}}$ are the gradient of the loss function $L(\cdot)$ with respect to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$, i.e., the $j$th row of $\bm{K}$ and $\bm{V}$ that correspond to the $j$th token. From (1) and (2), the weighted k-means clustering with gradients as weights is essentially to find a clustering strategy that minimizes the error caused by KV quantization. This also explains the effectiveness of the weighted k-means clustering used in this stage.
3.2 Online Anchor Token Selection
In the online stage, the gradient can no longer be used as AnS due to prohibitive memory and computational costs under real-time constraints. To this end, we here give the error propagation analysis of the self-attention (i.e., Attn). Specifically, the error propagation analysis is based on the perturbation bound of Attn that is related to each row of $\bm{K}$ and $\bm{V}$; see Theorem 3.1 (see Appendix B for proof).
Theorem 3.1.
Let $\delta\bm{K}$ and $\delta\bm{V}$ be the error perturbation terms corresponding to $\bm{K}$ and $\bm{V}$ respectively, and satisfy
$\|\delta\bm{K}\|_{L_{1}}\ll\|\bm{K}\|_{L_{1}}\quad\text{and}\quad\|\delta\bm{V}\|_{L_{1}}\ll\|\bm{V}\|_{L_{1}}.$
Then we have
$\|\texttt{Attn}(\bm{Q},\bm{K}+\delta\bm{K},\bm{V})-\texttt{Attn}(\bm{Q},\bm{K},\bm{V})\|_{L_{1}}$ (3)
$\lesssim$ $\sum_{j}\sum_{i}\bigl{\|}\bigl{(}\bm{V}^{\!\top}\mathrm{Diag}(\bm{A}_{i,:})(\bm{I}_{n}-\bm{e}\bm{A}_{i,:})\bigr{)}_{:,j}\bigr{\|}_{L_{1}}\|\bm{Q}_{i,:}\|_{L_{2}}\|\delta\bm{K}_{j,:}\|_{L_{1}}$
and
$\|\texttt{Attn}({\bm{Q}},{\bm{K}},\bm{V})-\texttt{Attn}({\bm{Q}},{\bm{K}},\bm{V}+\delta\bm{V})\|_{L_{1}}\leq\sum\limits_{j}\|\bm{A}_{:,j}\|_{L_{1}}\|\delta\bm{V}_{j,:}\|_{L_{1}},$ (4)
where $\bm{e}\in\mathbb{R}^{n}$ is a vector whose entries are all $1$.
We remark that the error propagation factors corresponding to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$ given in Theorem 3.1 can be regarded as the upper bound of the gradient of the self-attention operator related to $\bm{K}_{j,:}$ and $\bm{V}_{j,:}$.
The computation involving $\bm{K}$ in Theorem 3.1 entails substantial overhead, rendering it impractical for use during online inference. To address this, we propose a simplified variant that excludes the contribution of $\bm{V}$ to the quantization error of $\bm{K}$. By this simplification, we obtain the following reformulation. A detailed evaluation of the effectiveness of AnS is provided in Appendix C.
$\text{AnS}(\bm{K}_{j,:})=\sum_{i}\bm{A}_{i,j}(1-\bm{A}_{i,j})\cdot\left\|\bm{Q}_{i,:}\right\|_{2}$ (5)
During the prefill phase, AnS serves as an effective metric to identify anchor tokens that preserve critical contextual information by measuring their impact on attention outputs. In contrast, although AnS can still be computed during decoding, the anchor token it identifies may have already been quantized during prefill, preventing the preservation of its FP16 value for reducing accumulated errors. An important observation is that $\bm{A}$ and $\bm{A}\odot(\bm{E}-\bm{A})$ exhibit strong locality during decoding (see Appendix D), with anchor tokens primarily concentrated at the beginning and end of the sequence. In particular, the beginning tokens—referred to as sink tokens—are almost always identified as anchor tokens during prefill, indicating that our method effectively handle sink tokens. Motivated by this locality, we adopt a sliding window approximation of AnS during decoding.
3.3 Implementation
For the offline stage, the gradient of $\bm{K}$ and $\bm{V}$ is used as a weight to learn centroids. We implement it with a custom LinearWithAct to capture KV cache and corresponding gradients. Subsequently, we employ the weighted K-Means provided by cuML to perform efficient clustering.
For the online stage, AnS is derived from the error propagation factor given in Theorem 3.1 and (5). We design and implement a dedicated GPU kernel using Triton to efficiently compute it in conjunction with FlashAttention, enabling long-context inference. Since AnS requires reduction operations over the attention score matrix $\bm{A}$ and its transformed form $\bm{A}\odot(\bm{E}-\bm{A})$ along query (column-wise), directly fusing its computation into FlashAttention is impossible. To preserve the efficiency of FlashAttention, we decouple the AnS computation and execute it immediately afterward. To support this design, we extend FlashAttention to preserve three auxiliary tensors: the $L_{2}$ norm of each query vector, the key-wise (row-wise) sum, and the key-wise maximum of the matrix $\bm{Q}\bm{K}^{T}$. These additional outputs enable efficient reconstruction of the attention scores and allow AnS to be computed with minimal overhead. More details about the implementation are described in Algorithm 1 of Appendix E.
4 Experiments
In this section, we compare the performance of AnTKV with existing KV quantization methods through extensive experiments. The experimental setup, including models, benchmark datasets, evaluation metrics, and selected baselines, is as follows.
Models, Datasets, Metrics, and Parameter Settings. To validate the effectiveness of AnTKV in KV cache quantization and ensure the generality of results, we evaluate five representative models from the LLaMA and Mistral families: LLaMA-2-7B, LLaMA-3-8B, LLaMA-3-8B-Instruct, Mistral-7B, and Mistral-7B-Instruct-v0.3. We use 128 samples of length 2048 from the WikiText2-v1 training set as the calibration. Perplexity is evaluated on WikiText-2 and C4, while zero-shot accuracy on MMLU hendrycks2021mmlu , ARC-C clark2018arc , MathQA amini2019mathqa , and PIQA bisk2020piqa is used to assess understanding and reasoning. Long-context performance is measured using LongBench bai2024longbench . For perplexity and zero-shot benchmarks, quantized KV cache is employed for attention. In LongBench, the FP16 KV cache is first utilized to compute attention outputs, before quantization is applied.
[…]
4.1 Perplexity Results
Perplexity is a standard metric that is widely used to evaluate the quality of the output of LLMs, with lower values indicating better performance. The perplexity results for different KV quantization approaches on WikiText-2 and C4 are presented in Table 1. The results in this table indicate that the proposed AnTKV consistently achieves competitive or superior perplexity across various bit-widths and model architectures. Under 4-bit and 2-bit quantization, AnTKV achieves competitive or superior performance compared to existing approaches. In the 1-bit and sub-bit regimes, AnTKV significantly outperforms all baselines. It is worth mentioning that on the C4 dataset under sub-bit quantization, the baseline methods exhibit extremely high perplexity due to anchor tokens not being well captured by fixed centroids. Thanks to effective AnS design and innovative Anchor Token selection, AnTKV substantially reduces perplexity, e.g., from 66.28 to 14.42 (LLaMA-3-8B, 0.75-bit).
Table 1: The evaluations are conducted using each model’s maximum context length, specifically 4096 for LLaMA-2-7B and 8192 for LLaMA-3-8B and Mistral-7B. "Ours" refers to our method (AnTKV) without using any anchor tokens, while Ours-1% denotes AnTKV where 1% of the tokens are treated as anchor tokens and retained in FP16. For simplicity, the reported bit-widths do not consider the effect of centroids.
Bit LLaMA-2-7B LLaMA-3-8B Mistral-7B
Dataset WikiText2 C4 WikiText2 C4 WikiText2 C4
Baseline 16 5.12 6.63 5.54 7.10 4.73 5.66
RTN 4 5.66 7.31 7.89 8.79 7.34 5.91
SKVQ 5.16 6.67 5.64 7.19 4.97 5.68
KVQuant-1% 5.13 6.65 5.56 7.12 4.78 5.72
CQ 5.14 6.67 5.58 7.84 4.79 5.74
Ours 5.18 6.76 5.61 7.69 4.76 5.69
Ours-1% 5.15 6.68 5.59 7.16 4.74 5.67
RTN 2 4708 4708 2841 2113 573 477
SKVQ 5.54 7.21 6.73 8.31 5.21 6.14
KVQuant-1% 5.49 7.02 6.11 7.65 5.19 6.10
CQ 5.42 7.23 6.09 18.71 5.11 6.17
Ours 5.51 7.45 6.10 16.96 5.08 6.18
Ours-1% 5.34 7.02 5.97 7.68 4.95 5.97
SKVQ 1 12643 12819 108879 86426 3524 2741
KVQuant-1% 21.55 51.84 14.80 13.95 15.36 14.24
CQ 7.75 12.49 9.56 81.74 7.25 9.89
Ours 7.92 13.01 9.62 74.47 7.32 10.51
Ours-1% 6.50 9.40 8.51 12.51 6.32 8.44
CQ 0.75 8.39 14.32 11.18 72.05 7.64 11.72
Ours 8.21 14.27 10.41 66.28 7.41 11.72
Ours-1% 6.55 9.75 8.97 14.42 6.43 9.08
CQ 0.375 14.82 33.59 22.80 103.5 13.20 26.34
Ours 13.37 30.51 17.70 103.5 11.65 23.98
Ours-1% 8.75 15.86 13.41 34.08 8.87 14.87
4.2 Understanding and Reasoning Benchmark
To ensure that AnTKV exhibits broad understanding and reasoning capabilities, we evaluate its performance on four representative benchmarks: MMLU, ARC-Challenge, PIQA, and MathQA, using both LLaMA-8B-Instruct and Mistral-7B-Instruct. These benchmarks respectively assess multi-domain knowledge reasoning (MMLU), complex question answering (arc_challenge), commonsense reasoning (PIQA), and mathematical problem-solving (MathQA). Figure 4 illustrates the comparison results of AnTKV with SKVQ and CQ. Similarly, it can be seen that AnTKV consistently outperforms SKVQ and CQ across all four benchmarks under low-bit quantization, especially at 1-bit and sub-bit levels.
Figure 4: Understanding and Reasoning ability evaluation accuracy results on MMLU, ARC-C, PIQA, and MathQA under different quantization bit-widths.
4.3 Long-Context Benchmark
To validate the effectiveness of AnTKV in handling long-context, We conduct several experiments on the LLaMA-8B-Instruct model using the LongBench benchmark, a diverse collection of tasks such as question answering, retrieval, and summarization, designed to systematically evaluate long-context understanding in language models. We report results on eleven representative sub-tasks from LongBench, along with averaged performance. Due to alignment issues of KIVI and SKVQ, we exclude the triviaqa and gov_report sub-tasks from the comparison. As shown in Figure 5, AnTKV preserves nearly FP16 at 4- and 2-bit quantization across almost all tasks. At the 1-bit quantization, the performance of KIVI and SKVQ has a significant drop. In contrast, AnTKV and CQ still maintain a relatively high accuracy. To further investigate the robustness under aggressive compression, we compare AnTKV and CQ in both sub-bit levels. Figure 5 shows that AnTKV consistently outperforms CQ. Notably, when quantizing down to 0.375-bit, AnTKV maintains a tolerable performance degradation, with the average score dropping from 46.2 to 34.5.
Figure 5: The evaluation accuracy results on LongBench under different KV cache quantization bit-widths.
4.4 Efficiency
In this experiment, we evaluate the efficiency of our AnTKV implementation against the FlashAttention-accelerated FP16 baseline on LLaMA-3-8B using a single A100-80GB GPU. As shown in Figure 6, AnTKV substantially extends the maximum context length from 128K to 384K. In long-context scenarios, we observe that intermediate activations consume a significant portion of memory. By applying a series of in-place operations, AnTKV ultimately supports up to 810K tokens under 1-bit quantization and 840K under 0.375-bit quantization. More importantly, AnTKV sustains low memory consumption while maintaining acceptable latency. Under 1-bit quantization, a noticeable increase in prefill latency is observed only at the 128K context length, which represents the upper limit supported by most current models. It is observed that 0.375-bit quantization introduces higher latency than its 1-bit. This increase in latency arises from the use of a larger number of centroids, which are employed to enhance model accuracy.
To evaluate decoding efficiency, we evaluate the throughput of AnTKV under a fixed context length of 1K tokens. As shown in Figure 7, AnTKV supports significantly larger batch sizes compared to the baseline method. By substantially reducing KV cache access, it improves throughput across all bit-widths. Particularly, the maximum throughput reaches 3.5x that of the baseline method under 1-bit quantization.
Figure 6: Memory and latency comparison. Bars show memory usage (red: 1-bit, green: 0.375-bit); lines indicate prefill latency across context lengths.
Figure 7: Decoding throughput comparison. Our method supports larger batch sizes and achieves higher throughput.
4.5 Ablation Study
We conduct a series of experiments to answer the following questions.
Q1: How does the model performance change as the number of anchor tokens increases?
As the number of anchor tokens increases, the performance loss decreases rapidly at first, as shown in Table 1. However, the marginal benefit diminishes with more anchor tokens. Detailed results are provided in Appendix F.
Q2: Does the calibration set affect the performance?
We find that for VQ-based methods, the calibration set does have some impact on performance under low-bit settings. However, as shown in Table 1 (LLaMA-3-8B, 1-bit, C4), retaining anchor tokens effectively mitigates the performance drop caused by calibration set variation. More detailed results can be found in Appendix G.
5 Limitation & Conclusion
Although AnTKV demonstrates its advantages in experiments, it also has several limitations. First, more accurate AnS for tokens and higher performance implementations for its computation may be possible. AnTKV demonstrates strong potential in LLM serving by substantially reducing the size of the KV cache, which in turn alleviates I/O and memory constraints to a significant extent. Nevertheless, further empirical validation is required.
We observed from several experiments that different tokens exhibit varying sensitivity to KV quantization, which opens a new opportunity for a better memory-accuracy trade-off. To this end, we present AnTKV, a two-stage token-aware VQ framework that involves offline gradient weighted centroid learning and online anchor token selection guided by AnS. Moreover, to ensure the efficiency of AnTKV in LLM inference, we implement a lightweight GPU kernel compatible with FlashAttention to compute AnS with minimal overhead. Extensive experiments show that AnTKV achieves superior performance under low and ultra-low-bit quantization, significantly extending the maximum context length (up to 840K tokens) on a single A100 GPU while maintaining low perplexity and high throughput. Our results demonstrate the potential of token-aware quantization in scaling LLM inference to longer contexts without sacrificing accuracy.
Appendix A Distribution of pre- and post- RoPE Key
To identify a quantization strategy better suited for $\bm{K}$ vertor quantization, we compare the distribution of $\bm{K}$ before and after applying RoPE. Figure 8 presents a visualization of the pre- and post- RoPE $\bm{K}$. We observe that, compared to the post-RoPE $\bm{K}$, the pre-RoPE $\bm{K}$ exhibits smaller inter-cluster distances and lower intra-cluster variance, which contributes to reduced quantization error.
Figure 8: Distribution of pre- and post- RoPE Key. We sampled a 2048-length sentence from WikiText2 and generated pre- and post- RoPE Key on the LLaMA-3-8B model.
Appendix B Proof of Theorem 3.1
For $\bm{K}$, we have
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}=\|\left(\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}\right)-\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}^{T}\widetilde{\bm{K}}}{\sqrt{d}}\right)\right)\bm{V}\|_{L_{1}}.\end{split}$ (6)
The key to estimating the bound of (6) lies in the analysis of
$\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}\right)-\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right),$ (7)
whose $(i,j)$th entry is represented as
$\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}-\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}.$ (8)
Since $\|\delta\bm{K}\|_{L_{1}}\ll\|\bm{K}\|_{L_{1}}$, and by the first-order approximation $\exp{(x+\delta x)}\approx\exp{(x)}(1+\delta x)$, (8) can be approximated as
$\begin{split}\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\right)-\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)\right)}{\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\right)\left(\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)\right)}\\ =\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}\left(1+\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}\\ \approx\frac{\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{j,:}^{T}}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}_{s,:}^{T}}{\sqrt{d}}\right)}}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}\\ =\bm{A}_{i,j}\cdot\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}.\end{split}$
For convenience, we denote $\frac{\widetilde{\bm{Q}}\widetilde{\delta\bm{K}}^{T}}{\sqrt{d}}$ and $\left[\frac{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)\left(\frac{\widetilde{\bm{Q}}_{i,:}\left(\delta\bm{K}_{j,:}^{T}-\delta\bm{K}_{s,:}^{T}\right)}{\sqrt{d}}\right)}}{\sum\limits_{s}\exp{\left(\frac{\widetilde{\bm{Q}}_{i,:}\widetilde{\bm{K}}^{T}_{s,:}}{\sqrt{d}}\right)}}\right]_{n\times n}$ as $\bm{X}$ and $\bm{Y}$ respectively. Then (7) can be approximated as $\left(\bm{A}\odot\bm{Y}\right)\bm{V}$, and by the property of Kronecker product, we have
$\begin{split}\texttt{Vec}\left((\bm{A}\odot\bm{Y})\bm{V}\right)=\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Vec}\left(\bm{A}\odot\bm{Y}\right)=\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y}).\end{split}$
Further, we can obtain
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K}+\delta\bm{K},\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&\approx\|\left(\bm{I}_{n}\otimes\bm{V}^{T}\right)\texttt{Diag}\left(\texttt{Vec}(\bm{A})\right)\texttt{Vec}(\bm{Y})\|_{L_{1}}\\ &=\sum\limits_{i}\|\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\bm{X}_{i,:}^{T}\|_{L_{1}}\\ &\leq\sum\limits_{i}\sum\limits_{j}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}|\widetilde{\bm{Q}}_{i,:}\widetilde{\delta\bm{K}}_{j,:}^{T}|\\ &\leq\sum\limits_{j}\sum\limits_{i}\|\left(\bm{V}^{T}\texttt{Diag}\left(\bm{A}_{i,:}\right)\left(\bm{I}_{n}-\bm{e}\bm{A}_{i,:}\right)\right)_{:,j}\|_{L_{1}}\|{\bm{Q}}_{i,:}\|_{2}\|{\delta\bm{K}}_{j,:}\|_{L_{1}}.\\ \end{split}$ (9)
For $\bm{V}$, we have
$\begin{split}\|\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}+\delta\bm{V}\right)-\texttt{Attn}\left(\bm{Q},\bm{K},\bm{V}\right)\|_{L_{1}}&=\|\texttt{Softmax}\left(\frac{\widetilde{\bm{Q}}\widetilde{\bm{K}}^{T}}{\sqrt{d}}\right)\delta\bm{V}\|_{L_{1}}\\ &=\|\bm{A}\delta\bm{V}\|_{L_{1}}=\sum\limits_{i,k}|\sum\limits_{j}\bm{A}_{i,j}\delta\bm{V}_{j,k}|\\ &\leq\sum\limits_{i,k}\sum\limits_{j}\bm{A}_{i,j}|\delta\bm{V}_{j,k}|\\ &=\sum\limits_{j}\left(\sum\limits_{i}\bm{A}_{i,j}\right)\left(\sum\limits_{k}|\delta\bm{V}_{j,k}|\right)\\ &=\sum\limits_{j}\|\bm{A}_{:,j}\|_{L_{1}}\|\delta\bm{V}_{j,:}\|_{L_{1}}.\end{split}$ (10)
Appendix C The Effectiveness of AnS
We quantized each token across layers and heads, and separately recorded the errors in the attention outputs. As shown in Figure 9 and (5), it can be observed that the relative values derived from our AnS, as defined in Theorem 3.1, closely align with the actual outputs of the attention. For a fair comparison, the output errors for $\bm{K}$ also exclude the contribution of $\bm{V}$.
Figure 9: The effectiveness of AnS.
Appendix D AnS Distribution During Decoding
To illustrate the distribution of AnS during decoding, we sampled prompts from Qasper within LongBench for visualization. As shown in Figure 10, we present the distribution of AnS($\bm{K}$) and AnS($\bm{V}$) across different layers and heads, specifically when decoding the first token. Our observations reveal that high AnS values during decoding are predominantly concentrated on adjacent tokens and at the attention sink tokens. Since sink tokens often lead to significant error propagation and can be dynamically identified by AnS during prefill, we simplify the design of AnS during decoding by employing a sliding window to ensure model performance.
Figure 10: AnS Distribution on Sampled Prompts from Qasper Using LLaMA-3-8B-Instruct During First-Token Decoding.
Appendix E Computational Procedure of AnS in the Online Stage
Algorithm 1 The computation of AnS in the online stage.
1:  Input: Query ($\bm{Q}$), key ($\bm{K}$), value ($\bm{V}$)
2:  Output: AnS of KV, i.e., $\text{AnS}({\bm{K}})$ and $\text{AnS}({\bm{V}})$
3:  $(\bm{O},\bm{L},\bm{M},\left\|\bm{Q}_{i,:}\right\|_{L_{2}})\leftarrow\texttt{FlashAttention}(\bm{Q},\bm{K},\bm{V})$
4:  for each block key index $j$ in parallel (assigned to GPU block) do
5:     for each block query index $i$ do
6:        $\bm{S}_{i,j}\leftarrow\langle\bm{Q}_{h,i},\bm{K}_{h,j}\rangle$
7:        $\bm{A}_{i,j}\leftarrow\exp(\bm{S}_{i,j}-\bm{M}_{h,i})/\bm{L}_{h,i}$
[…]
10:     end for
11:  end for
12:  return $\text{AnS}({\bm{K}})$, $\text{AnS}({\bm{V}})$
Appendix F Calibration Set Impact
As shown in the Table 2, we observe that for VQ-based quantization in the ultra-low-bit regime, the calibration set significantly impacts the perplexity results. However, AnTKV with $1\%$ anchor tokens not only substantially reduces the PPL but also greatly mitigates the effect of different calibration sets.
Table 2: Perplexity experiment results on Mistral-7B, using the W2 and C4 training sets respectively as Calibration Sets. The "Vset" is the validation set related to W2 and C4. "Calib Set" represents "Calibration Set".
Bits Vset 4 2 1 0.75 0.375
Calib Set W2 C4 W2 C4 W2 C4 W2 C4 W2 C4
Ours W2 4.76 5.69 5.08 6.18 7.32 10.51 7.32 10.51 11.65 23.98
C4 4.79 5.69 5.32 6.15 10.14 10.09 10.90 10.80 24.16 19.95
Ours-1% W2 4.74 5.67 4.95 5.97 6.32 8.44 6.32 8.44 8.87 14.87
C4 4.75 5.66 5.02 5.94 7.13 8.13 7.79 8.56 12.87 13.07
To further investigate the impact of the calibration set on model performance, we used C4 as a calibration set to evaluate several subtasks within LongBench (qasper, trec, samsum, lcc, ropebench-p). As shown in the Table 3, we observed that there were some differences in the results of Trec and Repobench-p when using Wikitext-2 and C4, while the differences were not significant for the other tasks.
Table 3: Performance on LongBench Subtasks with Wikitext-2 (W2) and C4 Calibration Sets at Different Bits using LLaMA-3-8B-Instruct. "Calib Set" represents "Calibration Set", and "repobc-p" represents "repobench-p".
Bits 4 2 1 0.75 0.375
Calib Set W2 C4 W2 C4 W2 C4 W2 C4 W2 C4
qasper 40.46 39.98 39.04 38.2 25.95 26.51 25.48 25.49 22.41 23.27
trec 69.33 69.33 67 64.67 38.67 42.33 39.67 41.33 38 38
samsum 40.2 40.27 38.61 38.22 30.0 30.3 29.57 29.29 25.5 24.82
lcc 59.84 59.07 60.94 59.15 53.97 53,93 52.97 52.01 49.61 49.79
repobc-p 44.24 41.3 45.29 42.68 38.53 37.94 37.87 38.02 34.54 34.71
Appendix G Anchor Tokens Number Impact
To investigate the impact of the number of anchor tokens on model performance, we conducted Perplexity evaluations on Mistral-7B and LLaMA-3-8B, both with a context length of 8192. We performed evaluations using no anchor tokens and with anchor token percentages of 1% (82), 2% (164), 5% (410), 10% (820), 15% (1230), and 20% (1640). The corresponding results are presented in Figures 11 and 12. For the 2-bit and 4-bit results, using $1\%$ of anchor tokens kept the error within 0.6 compared to FP16. However, for the 1-bit and sub-bit results, we needed to increase the number of anchor tokens to control the error within an acceptable range. Nevertheless, AnTKV provides a feasible technical pathway for ultra-low-bit quantization of the KV cache.
Figure 11: Perplexity results on LLaMA-3-8B with varying anchor token numbers.
Figure 12: Perplexity results on Mistral-7B with varying anchor token numbers.
To further investigate the impact of anchor token numbers on downstream tasks, we evaluated different anchor token numbers on the Trec and Qasper subtasks of LongBench under ultra-low-bit quantization settings. For convenience, we approximated $1\%$ of the anchor token number as 64. The results are shown in Figure 13. The results indicate that Trec’s performance increases linearly with the number of anchor tokens, eventually reaching a result within 10 points of the baseline when using $10\%$ anchor tokens. In contrast to Trec, Qasper’s accuracy remains stable with a small number of anchor tokens but shows a significant performance improvement when increased to $10\%$. We believe this difference stems from the inherent properties of each task.
Figure 13: Trec and Qasper results on LLaMA-3-8B-Instruct with varying anchor token numbers.
Appendix H Models, Datasets and Code
H.1 Models Checkpoints
H.2 Datasets
H.3 Code
Our code and the trained centroids will be publicly released upon the acceptance of this paper.
Appendix I Broader Impact
In this work, we observe that tokens in the KV cache exhibit heterogeneous sensitivity to quantization, an insight that challenges the prevailing assumption of uniform importance across tokens. This observation paves the way for a new paradigm of importance-aware compression in LLMs. Our proposed method, AnTKV, embodies this paradigm through a two-stage, token-aware vector quantization framework. By selectively preserving high-impact tokens, AnTKV achieves strong performance even under ultra-low-bit quantization. Given its effectiveness and compatibility with high-throughput inference, AnTKV holds the potential to be widely adopted in future LLM compression and deployment practices.